OpenStudy (anonymous):

The measure of each exterior angle of a regular polygon is 30°. How many sides does the polygon have? Answer 10 12 14 15

OpenStudy (anonymous):

i thank its 12 but im not sure

OpenStudy (anonymous):

We'll, let's try this one out. 150n=180(n-2); 150n=180n-360, 30n=360, n=12. Do you understand?